OpenStudy (aravindg):

let f(x)= sin x ,g(x)=2x and h(x)=cos x,prove that fog =go(fh)

OpenStudy (anonymous):

it is go(f*h)=go(sinx*cosx)=g(sinx*cosx)=2sinxcosx=sin2x

OpenStudy (anonymous):

fog=f(2x)=sin2x got it same i think

OpenStudy (aravindg):

wow site has changed

OpenStudy (amistre64):

sirir got it i believe

OpenStudy (anonymous):

f(gx)=sin(2x) g(f*h)=2(sinx*cosx)

OpenStudy (anonymous):

yh amistre is right srir got it

OpenStudy (amistre64):

f(g(x)) = sin(g(x)) = sin(2x) g(f(x)*h(x)) = 2(f(x)*h(x)) = 2(sin(x) * cos(x)) = 2sin(x)cos(x) do we have to prove the identity: sin(2x) = 2sin(x)cos(x) as well

OpenStudy (anonymous):

^thts the only way this question would be harder, really though i wonder what grade u are in ARavind

OpenStudy (aravindg):

11th